OpenStudy (anonymous):

Help me differentiate y=x^4 -2x^2+8x

OpenStudy (anonymous):

ake ln of both sides to get ln y = ln(x^8x) use log rules to move the 8x ln y = 8x*ln x take the derivative of both sides (using chain rule on ln y since y is a function of x) y' / y = (8)(ln x) + (8x)(1/x) y' / y = 8(ln x) + 8 multiply both sides by y y' = y [8(ln x)+8] substitute x^8x for y y' = x^8x[8(ln x)+8]

OpenStudy (micahm):

move this to math thanks

OpenStudy (micahm):

this is English your in

OpenStudy (micahm):

close this and move it to math and people will help you there

OpenStudy (anonymous):

stop talking

OpenStudy (micahm):

cus